# Code Simulation as a Proxy for High-order Tasks in Large Language Models

### v2

@@ 1 Introduction
A major area of interest at the time of writing is understanding the capabilities of Large Language Models (LLMs) beyond the tasks of language understanding and generation. Many benchmarks, including Theory of Mind (Rabinowitz et al., 2018), planning (Hao et al., 2023; Ouyang et al., 2022), and high-order reasoning (Webb et al., 2023), necessitate turning a prompt, expressed in natural language, into a procedure that must then be carried on faultlessly. Consider a problem where two agents interact and exchange goods, as in Figure 1. An LLM prompted to compute the number of goods at the end of an iteration should be able to sum and assign variables correctly. Such a naturalistic problem has an intrinsic algorithmic nature as code (Figure 1, centre). This work centres on this analogy and turns it into an experimental pipeline to assess the capabilities of LLMs on reasoning tasks via equivalent code simulation.
@@
Figure 1: Left: an example of the naturalistic vs. synthetic good exchange settings. The former describes, in natural language, two agents who exchange goods; the latter is an equivalent formulation in code. While GPT-3.5-Turbo performs better on the synthetic task (a “simulation gap”), performance in the synthetic and naturalistic tasks are strongly correlated with respect to the control variable, i.e., the number of operations/exchanges. We conduct experiments on $30$ samples per instruction class with $\{10,20,30,40,50\}$ interactions/ lines of code.
This paper shows that code simulation is a scalable proxy for assessing some core reasoning capabilities of LLMs; this is in contrast with generating naturalistic datasets, which is resource- and time-expensive. While not all the reasoning tasks can be equivalently formulated as code, we showcase the flexibility of our framework by pairing five non-trivial naturalistic tasks with their coding counterpart and show that the performances of GPT-4, GPT-4o (OpenAI, 2023), and Llama3.1-405B (Grattafiori et al., 2024) correlate on them. As expected, our code simulation approach enables us to identify failure cases for step-by-step execution, namely memorisation and “lazy” pattern recognition. For these cases, we develop a minimal extension of Chain of Thought (Wei et al., 2022b) that mitigates these issues and allows benchmarking in the presence of memorisation. To further lay the groundwork for code simulation as a proxy of naturalistic reasoning tasks, we conduct extensive experiments on pure code simulation and simple programming constructs on a variety of models, including GPT-3.5-Turbo, GPT-4, Jurassic2-Ultra, Llama-2-70B, Llama-3-70B and CodeLLaMA-34b-Instruct. The code to replicate the experiments in the main paper is available here, while the code to replicate the experiments on pure code simulation is available here. Click here to access the project’s website.
@@ 3 Methodology
Formally, a Language Model is defined as a function that predicts the next token (out of a finite vocabulary) conditioned on the sequence of previously fed/generated tokens, namely $\psi:V^{*}\xrightarrow{}\mathbb{P}(V)$. In our setting, a problem is specified as a tuple $(x,p)$, where $x$ instructs the model to solve a problem $p$ expressed either as code or as an equivalent naturalistic task. Both settings express the same question, with the ground truth label obtained by running the code with an interpreter/compiler $\Omega$ and compared, for correctness, against the model’s answer. We select Python 3 as our programming language for coding problems as it represents the language of reference in most LLMs such as Code-LLaMA (Rozière et al., 2023) and is among the most covered languages in open-source LLMs (Gao et al., 2020; Scao et al., 2022) and network Q&A platforms such as Stack Exchange. Furthermore, its syntax, for simple programs, resembles that of pseudo-code, thus abstracting from complex programming constructs.
@@ 3.1 Benchmarks
Another aspect of reasoning that complements the “extraneous cognitive load” is that of “intrinsic load” as the complexity of processing new information (Sweller & Chandler, 1991). One can make a Straight line and Good exchange problem more complex by requiring to keep track of multiple variables/objects simultaneously. A model has to store, access, and modify all the variables consistently to return the correct result. The Clique good exchange is a Multi-Agent Scenario that, paired with the Parallel paths, tests this setting. Figure 2 (left) illustrates an example of such a setting. Similarly to the previous benchmarks, in Appendix B.4, we evaluate a richer set of operations, such as logical and/or, that are hard to model in naturalistic settings but can serve as the basis of increasingly complex synthetic benchmarks.
@@ 4.2 The “Simulation Gap”: Between Shortcuts and Memorisation
We study GPT-3.5-Turbo, GPT-4, and Llama-3-70B and run $3$ independent runs of $30$ experiments each are shown in Figure 5. We analyse GPT-3.5-Turbo here as it results in the most interesting findings, though the same trends affect GPT-4 and Llama-3; see Appendix B.6. Beyond confirming the implicit bias of Transformers towards ordered sequences (Dufter et al., 2022) (i.e., they tend to output ordered sequences), Figure 5 (left) shows that LLMs often provide the correct result for classic sorting algorithms such as Insertion Sort. However, with less common algorithms, LLMs trace the expected execution but fail even for vectors of few inputs. In summary, we find that (i) standard sorting algorithms (e.g., Insertion and Quick Sort) lead to more accurate results; (ii) there is a weak correlation between algorithm complexity and the accuracy of its simulation, thus reinforcing the hypothesis that in this case, the model is not simulating the procedure; (iii) there is a weak correlation between the length of the input vector and the accuracy of a model.
@@
As the main research question of this work is whether one can use code as a proxy of high-order/reasoning tasks, it is important to investigate the role of memorisation on common algorithms, as they may bias the comparison to naturalistic, non-memorised tasks. We paired five standard sorting algorithms with slight variations that neither affect the code length nor their computational complexity. Yet, their semantics are slightly changed to produce a different output. We also introduce a prompting technique that, on top of CoT prompting, explicitly instructs a model to simulate a routine step by step: we name this prompting method Chain of Simulation, or CoSm (the prompt we use is reported in Appendix C). We investigated the following algorithms and their variations: (i) the Fibonacci sequence paired with Padovan, a slight variation that modifies the return condition; (ii) ascending Bubble Sort, paired with the descending routine; (iii) the Gauss algorithm, paired with a variation that, instead of summing the first $n$ natural numbers, adds even and subtracts odd numbers; (iv) a primality test paired with the same routine on the successor of the input; (v) and the sum of the first $n$ Collatz numbers, paired with a variation that returns the sum of the even numbers in a Collatz sequence. All the functions and their variations were first anonymised to avoid bias towards known function names and implemented with the code that appears most frequently on GitHub. We report their implementation in Appendix D.2. As shown in Figure 6 (top), GPT-3.5-Turbo, GPT-4 and Llama-3-70B are accurate on each classic algorithm, but their accuracy dropped significantly with the variations. On the other hand, their accuracy is marginally improved when we prompt them explicitly to simulate the routine with CoSm (Figure 6, bottom). We observe that LLMs anticipate the behaviour of a function by looking at specific dominating patterns. As shown in Figure 7 (right), GPT-3.5-Instruct completes a template of the Fibonacci function with tokens compatible with the function’s scope. On Padovan, there is a non-negligible probability (the third most likely token, in red) that the predicted function is Fibonacci. These results question the induction head mechanism (Olsson et al., 2022), as a slight variation of a task where an LLM is accurate results in errors of large order of magnitude. Furthermore, this particular type of uncertainty goes unnoticed with methods that detect memorisation (Shi et al., 2023), as shown in Figure 7 (right). While anti-memorisation techniques are effective with textual inputs (e.g., the Wikipedia dataset), they fail on code. We hypothesise that low-frequency tokens are informative for natural language but not on code as it is more structured, with low-likelihood tokens (in magenta) that are not predictive of a model’s memorisation.
@@ 5 Conclusions and Future Work
In this work, we introduce a way to evaluate some high-order capabilities of LLMs via synthetic tasks in the form of code. We show that many reasoning tasks have a natural reformulation as code, which is easy to obtain and scale. Our experiments show the feasibility of our approach and pave the way to synthetic benchmarks: on the other hand, code simulation may suffers from memorisation (for common routines such as sorting), and one has to carefully consider these drawbacks when designing synthetic task as a proxy of high-order reasoning.

### v3

@@ 1 Introduction
A major area of interest at the time of writing is understanding the capabilities of Large Language Models (LLMs) beyond the tasks of language understanding and generation. Many benchmarks, including Theory of Mind (Rabinowitz et al., 2018), planning (Hao et al., 2023; Ouyang et al., 2022), and high-order reasoning (Webb et al., 2023), necessitate turning a prompt, expressed in natural language, into a procedure that must then be carried out faultlessly. Consider a problem where two agents interact and exchange goods, as in Figure 1. An LLM prompted to compute the number of goods at the end of an iteration should be able to sum and assign variables correctly. Such a naturalistic problem has an intrinsic algorithmic nature as code (Figure 1, centre). This work centres on this analogy and turns it into an experimental pipeline to assess the capabilities of LLMs on reasoning tasks via equivalent code simulation.
@@
Figure 1: Left: an example of the naturalistic vs. synthetic good exchange settings. The former describes, in natural language, two agents who exchange goods; the latter is an equivalent formulation in code. While GPT-3.5-Turbo performs better on the synthetic task (a “simulation gap”), performance in the synthetic and naturalistic tasks is strongly correlated with respect to the control variable, i.e., the number of operations/exchanges. We conduct experiments on $30$ samples per instruction class with $\{10,20,30,40,50\}$ interactions/ lines of code.
This paper shows that code simulation is a scalable proxy for assessing some core reasoning capabilities of LLMs; this is in contrast with generating naturalistic datasets, which is resource- and time-expensive. While not all the reasoning tasks can be equivalently formulated as code, we showcase the flexibility of our framework by pairing five non-trivial naturalistic tasks with their coding counterpart and show that the performances of GPT-4, GPT-4o (OpenAI, 2023), and Llama3.1-405B (Grattafiori et al., 2024) correlate with them. As expected, our code simulation approach enables us to identify failure cases for step-by-step execution, namely memorisation and “lazy” pattern recognition. For these cases, we develop a minimal extension of Chain of Thought (Wei et al., 2022b) that mitigates these issues and allows benchmarking in the presence of memorisation. To further lay the groundwork for code simulation as a proxy of naturalistic reasoning tasks, we conduct extensive experiments on pure code simulation and simple programming constructs on a variety of models, including GPT-3.5-Turbo, GPT-4, Jurassic2-Ultra, Llama-2-70B, Llama-3-70B and CodeLLaMA-34b-Instruct. The code to replicate the experiments in the main paper is available here, while the code to replicate the experiments on pure code simulation is available here. Click here to access the project’s website.
@@ 3 Methodology
Formally, a Language Model is defined as a function that predicts the next token (out of a finite vocabulary) conditioned on the sequence of previously fed/generated tokens, namely $\psi:V^{*}\xrightarrow{}\mathbb{P}(V)$. In our setting, a problem is specified as a tuple $(x,p)$, where $x$ instructs the model to solve a problem $p$ expressed either as code or as an equivalent naturalistic task. Both settings express the same question, with the ground truth label obtained by running the code with an interpreter/compiler $\Omega$ and compared, for correctness, against the model’s answer. We select Python 3 as our programming language for coding problems as it represents the language of reference in most LLMs, such as Code-LLaMA (Rozière et al., 2023), and is among the most covered languages in open-source LLMs (Gao et al., 2020; Scao et al., 2022) and network Q&A platforms such as Stack Exchange. Furthermore, its syntax, for simple programs, resembles that of pseudo-code, thus abstracting from complex programming constructs.
@@ 3.1 Benchmarks
Another aspect of reasoning that complements the “extraneous cognitive load” is that of “intrinsic load” as the complexity of processing new information (Sweller & Chandler, 1991). One can make a Straight line and a Good exchange problem more complex by requiring one to keep track of multiple variables/objects simultaneously. A model has to store, access, and modify all the variables consistently to return the correct result. The Clique good exchange is a Multi-Agent Scenario that, paired with the Parallel paths, tests this setting. Figure 2 (left) illustrates an example of such a setting. Similarly to the previous benchmarks, in Appendix B.4, we evaluate a richer set of operations, such as logical and/or, that are hard to model in naturalistic settings but can serve as the basis of increasingly complex synthetic benchmarks.
@@ 4.2 The “Simulation Gap”: Between Shortcuts and Memorisation
We study GPT-3.5-Turbo, GPT-4, and Llama-3-70B and run $3$ independent runs of $30$ experiments each are shown in Figure 5. We analyse GPT-3.5-Turbo here as it results in the most interesting findings, though the same trends affect GPT-4 and Llama-3; see Appendix B.6. Beyond confirming the implicit bias of Transformers towards ordered sequences (Dufter et al., 2022) (i.e., they tend to output ordered sequences), Figure 5 (left) shows that LLMs often provide the correct result for classic sorting algorithms such as Insertion Sort. However, with less common algorithms, LLMs trace the expected execution but fail even for vectors of a few inputs. In summary, we find that (i) standard sorting algorithms (e.g., Insertion and Quick Sort) lead to more accurate results; (ii) there is a weak correlation between algorithm complexity and the accuracy of its simulation, thus reinforcing the hypothesis that in this case, the model is not simulating the procedure; (iii) there is a weak correlation between the length of the input vector and the accuracy of a model.
@@
As the main research question of this work is whether one can use code as a proxy of high-order/reasoning tasks, it is important to investigate the role of memorisation on common algorithms, as they may bias the comparison to naturalistic, non-memorised tasks. We paired five standard sorting algorithms with slight variations that neither affect the code length nor their computational complexity. Yet, their semantics are slightly changed to produce a different output. We also introduce a prompting technique that, on top of CoT prompting, explicitly instructs a model to simulate a routine step by step: we name this prompting method Chain of Simulation, or CoSm (the prompt we use is reported in Appendix C). We investigated the following algorithms and their variations: (i) the Fibonacci sequence paired with Padovan, a slight variation that modifies the return condition; (ii) ascending Bubble Sort, paired with the descending routine; (iii) the Gauss algorithm, paired with a variation that, instead of summing the first $n$ natural numbers, adds even and subtracts odd numbers; (iv) a primality test paired with the same routine on the successor of the input; (v) and the sum of the first $n$ Collatz numbers, paired with a variation that returns the sum of the even numbers in a Collatz sequence. All the functions and their variations were first anonymised to avoid bias towards known function names and implemented with the code that appears most frequently on GitHub. We report their implementation in Appendix D.2. As shown in Figure 6 (top), GPT-3.5-Turbo, GPT-4 and Llama-3-70B are accurate on each classic algorithm, but their accuracy dropped significantly with the variations. On the other hand, their accuracy is marginally improved when we prompt them explicitly to simulate the routine with CoSm (Figure 6, bottom). We observe that LLMs anticipate the behaviour of a function by looking at specific dominating patterns. As shown in Figure 7 (right), GPT-3.5-Instruct completes a template of the Fibonacci function with tokens compatible with the function’s scope. On Padovan, there is a non-negligible probability (the third most likely token, in red) that the predicted function is Fibonacci. These results question the induction head mechanism (Olsson et al., 2022), as a slight variation of a task where an LLM is accurate results in errors of a large order of magnitude. Furthermore, this particular type of uncertainty goes unnoticed with methods that detect memorisation (Shi et al., 2023), as shown in Figure 7 (right). While anti-memorisation techniques are effective with textual inputs (e.g., the Wikipedia dataset), they fail on code. We hypothesise that low-frequency tokens are informative for natural language but not on code as it is more structured, with low-likelihood tokens (in magenta) that are not predictive of a model’s memorisation.
@@ 5 Conclusions and Future Work
In this work, we introduce a way to evaluate some high-order capabilities of LLMs via synthetic tasks in the form of code. We show that many reasoning tasks have a natural reformulation as code, which is easy to obtain and scale. Our experiments show the feasibility of our approach and pave the way to synthetic benchmarks: on the other hand, code simulation may suffer from memorisation (for common routines such as sorting), and one has to carefully consider these drawbacks when designing a synthetic task as a proxy of high-order reasoning.
Acknowledgments
ELM is supported by the Alan Turing Institute. AGC is supported by the Economic and Social Research Council (ESRC) under grant ES/W003473/1, by the Fundamental Research priority area of The Alan Turing Institute, and by the Turing Defence and Security programme, through a partnership with the UK government in accordance with the framework agreement between GCHQ and the Alan Turing Institute. We thank Microsoft Research - Accelerating Foundation Models Research program for the provision of Azure resources to access OpenAI models.